Photochronograph Corporation (PC) manufactures time series photographic equipment. It is currently at its target debt-equity ratio of .8. It’s considering building a new $72 million manufacturing facility. This new plant is expected to generate aftertax cash flows of $8.4 million in perpetuity. The company raises all equity from outside financing. There are three financing options:

1.

A new issue of common stock: The flotation costs of the new common stock would be 6.7 percent of the amount raised. The required return on the company’s new equity is 15 percent.

2.

A new issue of 20-year bonds: The flotation costs of the new bonds would be 2.6 percent of the proceeds. If the company issues these new bonds at an annual coupon rate of 6 percent, they will sell at par.

3.

Increased use of accounts payable financing: Because this financing is part of the company’s ongoing daily business, it has no flotation costs, and the company assigns it a cost that is the same as the overall firm WACC. Management has a target ratio of accounts payable to long-term debt of .20. (Assume there is no difference between the pretax and aftertax accounts payable cost.)

What is the NPV of the new plant? Assume that PC has a 25 percent tax rate. (Do not round intermediate calculations and enter your answer in dollars, not millions, rounded to the nearest whole dollar amount, e.g., 1,234,567.)
Answer #1
Financing Option-1 Common stock
Floatation cost 6.70%
Required return 15%
Fund required 72000000
Which is net of floation cost hence to amount required to raise.
Total fund required= Fund after floation cost X 100
100-% of floation cost
= 72000000*100/100-6.7
= 7,71,70,418.01
Revised required return= Original Investment X Required rate of return
Investment after floatation cost
= (77170418.01*15%)/72000000*100
= 16.08
Present value of cash flow = 8400000/16.08 *100
5,22,38,805.97
Net present Value= present value of cash flow-Cost of building
= 52238805.9701493-72000000
= -1,97,61,194.03
Financing Option-2 Bond
Floatation cost 2.60%
Required return 6%
Fund required 72
Which is net of floation cost hence to amount required to raise.
Total fund required= Fund after floation cost X 100
100-% of floation cost
= 72000000*100/100-2.6
= 7,39,21,971.25
Revised required return= Original Investment X Required rate of return
Investment after floatation cost
= (73921971.25*6%)/72000000*100
Pretax Rate = 6.16
Post tax = 6.16*.75
= 4.62
Present value of cash flow Post Tax = 8400000/4.62 *100
= 136363636.4
Present value of cash flow Post Tax = 8400000/4.62 *100
= 181818181.8
Pre tax Net present Value= present value of cash flow-Cost of building
= 136363636.36-72000000
208363636
Post tax Net present Value= present value of cash flow-Cost of building
= 181818181.82-72000000
= 109818182
Financing Option-3 Accounts payable
Present Debt Equity ratio 0.8
Formula of debt equity = Total debt/ Equity
Project Cost 72000000
Total Financing = Debt +Equity
Total Financing = 1.8
Debt = 32000000
Equity = 40000000
Total = 72000000
Pre tax WaCC of Firm = (Equity weight * Rate of Return)+(Debt *Rate of interest)
Pre tax WaCC of Firm = ((1*.15)+(.8*.06))*100
= 19.8
Post Tax WaCC of Firm (Equity weight * Rate of Return)+(Debt *Rate of interest after tax)
= ((1*.15)+(.8*.06*.75))*100
= 18.6
Present value of Cash flow = 8400000/18.6 *100
45161290.32
Net present Value = 45161290.32-72000000
= -26838709.68